OpenStudy (anonymous):

Find the measurement of one angle in a 72-gon

OpenStudy (jdoe0001):

\(\large \begin{array}{rllll} a&\cdot &n=360\\ \uparrow &&\uparrow \\ \textit{external angle}&&\textit{number of sides}\\ \downarrow&&\downarrow\\ a&\cdot &72=360 \end{array}\) solve for "a" to get the external angle if you want the internal angle keep in mind that internal and external angles are both supplementary angles